# BS-tree: A gapped data-parallel B-tree

### v1

@@ 1. Introduction
The B+-tree is the dominant indexing method for DBMSs, due to its low and guaranteed cost for (equality and range) query processing and updates. It was originally proposed as a disk-based index, where the objective is to minimize the I/O cost of operations. As memories become larger and cheaper, main-memory and hardware-specific implementations of the B+-tree (Boyar and Larsen, 1992; Lehman and Carey, 1986; Rao and Ross, 1999, 2000; Bohannon et al., 2001; Kim et al., 2010; Fix et al., 2011; Levandoski et al., 2013; Kaczmarski, 2012; Shahvarani and Jacobsen, 2016; Yan et al., 2019; Kwon et al., 2023), as well as alternative access methods for in-memory data (Morrison, 1968; Askitis and Sinha, 2010; Böhm et al., 2011; Kissinger et al., 2012; Mao et al., 2012; Leis et al., 2013; Zhang et al., 2016; Binna et al., 2018; Zhang et al., 2018) have been proposed. The optimization objective in all these methods is minimizing the computational cost and cache misses during search. More recently, learned indices (Kraska et al., 2018; Galakatos et al., 2019; Ferragina and Vinciguerra, 2020; Ding et al., 2020a; Zhang and Gao, 2022; Wu et al., 2022; Kipf et al., 2020; Wu et al., 2021; Li et al., 2023; Zhang et al., 2024), which replace the inner nodes of the B+-tree by ML models have been suggested as a way for reducing the memory footprint of indexing and accelerating search at the same time.
@@
In this paper, we propose BS-tree, a B+-tree for main memory data, which is optimized for modern commodity hardware and data parallelism. BS-tree adopts the structure of the disk-based B+-tree (i.e., a balanced, multiway tree), setting the node size to a memory block that can be processed in parallel. At the heart of our proposal lies a data-parallel successor operator ($succ$), implemented using SIMD, which is applied at each tree level for branching during search and updates and for locating the search key position at the leaf level. To facilitate fast updates, without affecting SIMD-based search, we propose a novel implementation for gaps (unused positions) by duplicating keys. The main idea is that we write in each unused slot the next used key value in the node or a global MAXKEY value if all subsequent slots are unused. Our BS-tree construction algorithm initializes sparse leaf nodes with intentional gaps in them, in order to (i) delay possible splits and (ii) reduce data shifting at insertions. Splitting also adds gaps proactively. Finally, we apply a frame-of-reference (FOR) based compression method that allows nodes that use fixed-size memory blocks to have varying capacities, which saves space and increases data parallelism. The computational cost of search and update operations in BS-tree is $O(\log_{f}n)$, where $f$ is the capacity of the nodes, assuming that $f$ is selected such that each node can be processed by a (small) constant number of SIMD instructions.
@@
Novelty and contributions. There already exist several SIMD-based implementations of B-trees and k-ary search (Schlegel et al., 2009; Kim et al., 2010; Shahvarani and Jacobsen, 2016; Yan et al., 2019; Kwon et al., 2023; Graefe, 2024). In addition, updatable learned indices (Ding et al., 2020a) use gaps to facilitate fast updates. Finally, key compression in B-trees has also been studied in previous work (Bohannon et al., 2001). To our knowledge, our proposed BS-tree is the first B+-tree implementation that gracefully combines all these features, achieving at the same time minimal storage and high throughput. In particular, the use of duplicate key vales in unused gaps/slots allows (i) branchless, data-parallel SIMD search at each node, and (ii) efficient key insertions and deletions with limited shifting of keys within each node. In addition, our compression scheme allows for direct application of data-parallel search on compressed nodes. To our knowledge, using duplicate keys in gaps within each node for efficient data-parallel search and updates at the same time is novel and has not been supported by previous B-tree implementations (Graefe, 2024). We implement a version of optimistic lock coupling in BS-tree for concurrency control and extensively compare BS-tree with open-source single- and multi-threaded implementations of state-of-the-art non-learned and learned indices on widely used real datasets, to find that BS-tree and its compressed version consistently prevail in different workloads of reads and updates, typically achieving 1.5x-2x higher throughput than the best competitor.
@@ 2.1. B-tree
The B+-tree is considered the de-facto access method for relational data, having substantial advantages over hash-based indexing with respect to construction cost, support of range queries, sorted data access, and concurrency control (Graefe, 2011; Elmasri and Navathe, 1989; Graefe, 2024). As memory sizes grow, the interest has shifted to in-memory access methods (Zhang et al., 2015). Rao and Ross (Rao and Ross, 1999) were the first to consider the impact of cache misses in memory-based data structures; they proposed Cache-Sensitive Search Trees (CSS-trees), in which every node has the same size as the cache-line of the machine and does not need to keep pointers for the links between nodes, but offsets that can be calculated by arithmentic operations. Rao and Ross (Rao and Ross, 2000) also proposed the Cache Sensitive B+-tree (CSB+tree), which achieves cache performance close to CSS-Trees, while having the advantages of a B+-tree. Chen et al. (Chen et al., 2001, 2002) showed how prefetching can significantly improve the performance of index structures by reducing memory access latency. The pkB-tree (Bohannon et al., 2001) is an in-memory variant of the B-tree that uses partial-keys (fixed-size parts of keys), which reduce cache misses and improve search performance. Zhou and Ross (Zhou and Ross, 2003) investigated buffering techniques, based on fixed-size or variable-sized buffers, for memory index structures, aiming to avoid cache thrashing and to improve the performance of bulk lookup in relation to a sequence of single lookups. Graefe and Larson (Graefe and Larson, 2001) surveyed techniques that improve the perfomance of B+-tree by exploting CPU caches.
@@
The advent of SIMD instructions and GPUs opened new perspectives for in-memory index structures. In an early work, Zhou and Ross (Zhou and Ross, 2002) explored the use of SIMD to parallelize key database operations (such as scans, joins, and filtering), minimizing branch mispredictions. Schlegel et al. (Schlegel et al., 2009) present methods for SIMD-based k-ary search (find which out of k partitions contains a search key). FAST (Kim et al., 2010) optimizes k-way tree search by leveraging architecture-specific features such as cache locality, SIMD parallelism on CPUs, and massive parallelism on GPUs. (Fix et al., 2011) introduced a “braided” B+-tree structure optimized for parallel searches on GPUs, enabling lock-free traversal using additional pointers. Kaczmarski (Kaczmarski, 2012) proposed a bottom-up B+-tree construction and maintenance technique using CPU and GPU for bulk-loading and updates. Bw-Tree (Levandoski et al., 2013) is a highly scalable and latch-free B+-tree variant optimized for modern hardware platforms, including multi-core processors and flash storage. Hybrid B+-tree (Shahvarani and Jacobsen, 2016) leverages both CPU and GPU resources to optimize in-memory indexing, by dynamically balancing the workload between the CPU and the GPU. Yan et al. (Yan et al., 2019) proposed a B+-tree tailored for GPU and SIMD architectures. This structure decouples the “key region”, which contains keys of the B+-tree with the “child region”, which is organized as a prefix-sum array and stores only each node’s first child index in the key region. Kwon et al. (Kwon et al., 2023) proposed DB+-tree, a B+-tree with partial keys, that utilizes SIMD and other sequential instructions for fast branching. PALM (Sewall et al., 2011) is a parallel latch-free variant of the B+-tree, which is optimized for multi-core processors, enabling concurrent search and update operations. Other works explore the implementation of B-trees on flash memory (Wu et al., 2007; Agrawal et al., 2009; Li et al., 2010; Jørgensen et al., 2011; Na et al., 2012; Athanassoulis and Ailamaki, 2014; Jin et al., 2016; Wang et al., 2020), non-volatile memory (Chen and Jin, 2015; Liu et al., 2020) and hardware transactional memory (Siakavaras et al., 2020).
@@ 2.3. Other in-memory access methods
Besides B-trees, other data structures have also been used for in-memory indexing, especially trie-based ones, such as the HAT-trie (Askitis and Sinha, 2007, 2010), the generalized prefix tree (trie) (Böhm et al., 2011), KISS-TREE (Kissinger et al., 2012), and Masstree (Mao et al., 2012). Leis et al. (Leis et al., 2013) proposed a fast and space-efficient in-memory trie called ART, which dynamically adjusts its node sizes providing a compact and cache-efficient representation. ART uses lazy expansion and path compression to improve space utilization and search performance. Leis et al. proposed two synchronization protocols for ART in (Leis et al., 2016), which have good scalability despite relying on locks: optimistic lock coupling and the read-optimized write exclusion (ROWEX) protocol. Height Optimized Trie (HOT) (Binna et al., 2018) is an in-memory trie-based index that reduces tree height through path compression and node merging. SuRF (Succinct Range Filter) (Zhang et al., 2018) leverages succinct tries to provide a space-efficient solution for range query filtering.
@@ 2.4. Learned Indexing
The advent of fast and accurate machine learning techniques inspired the design of a new type of index structure, called learned index (Kraska et al., 2018). The main idea is to learn a cumulative distribution function (CDF) of the keys and define a Recursive Model Index (RMI) that replaces the inner nodes of a B+-tree by a hierarchy of models that can predict very fast the position of the search key. FITing tree (Galakatos et al., 2019) and PGM-index (Ferragina and Vinciguerra, 2020) build upon RMI with a focus on improving model performance. with provable worst-case bounds on query time and space usage. The RadixSpline (RS) (Kipf et al., 2020) learned index can be constructed in a single traversal of sorted data. ALEX (Ding et al., 2020a) is an updatable learned index structure, based on RMI. ALEX utilizes a gapped array layout that gracefully distributes extra space between elements based on the model’s predictions, enabling faster updates and lookups by exponential search. Other updatable learned indices include CARMI (Zhang and Gao, 2022), NFL (Wu et al., 2022), LIPP (Wu et al., 2021), and DILI (Li et al., 2023), and Hyper (Zhang et al., 2024). Refs. (Marcus et al., 2020; Kipf et al., 2019; Wongkham et al., 2022) provide comprehensive evaluations on updatable learned indices and traditional indices including many important findings, based on tests on several real-world datasets.
@@ 3.1. The structure
BS-tree follows the structure of the B+-tree. Each internal node of the tree fits up to $N$ references to nodes at the lower level and up to $N-1$ keys. Leaf nodes contain rid-key pairs, where a record-id (rid) is the address (potentially on the disk) of the record that has the corresponding key value. We assume that keys are unique (if not, rid is replaced by a pointer to a block that keeps the rid’s of all records having the corresponding search key). The storage of the rid’s is decoupled from the storage of the keys, i.e., they are stored in two different arrays, such that the rid array is accessed only if necessary (i.e., only if the key is found and we need access to the corresponding record). Similarly, the storage of keys in an internal node is decoupled from the storage of children addresses, to facilitate fast search, as we explain later. Each leaf node hosts the address of the next leaf in the total key order; Leaf chaining efficiently supports range queries, where query results are in consecutive tree leaves. For the BS-tree nodes we use a value of $N$ that facilitates fast and parallel search, as we will explain in Sections 5 and 6. For the efficient handling of updates, we allow gaps (i.e., unused slots) in nodes, similarly to previous work (Ding et al., 2020a; Wu et al., 2021; Li et al., 2023), as will be discussed in Section 4.
Figure 1 shows an example of a BS-tree, where each node holds up to $N-1=4$ keys. Each non-leaf node is shown as an array of $N$ node pointers (bottom) and $N-1$ keys (top) that work as separators. All keys in the subtree pointed by the $i$-th pointer are strictly smaller than the $i$-th key and greater than or equal to the $(i-1)$-th key (if $i>0$). Any unused key slots at the end of each node carry a special MAXKEY value (denoted by $\infty$ in the figure), which is larger than the maximum possible value in the key domain. For example, if keys are unsigned 64-bit integers, MAXKEY $=2^{64}-1$ and key values range in $[0,2^{64}-2]$.
@@ 3.2. Search within a BS-tree node
1 int succG(Node *v, uint64 skey) {
@@
2 int succG_SIMD(Node *v, uint64 skey) {
@@
Similarly, $succ_{\geq k}(v)=|\{x:x\in v.keys\land k>x\}|$ and the same code snippets can be used, by replacing comparison >= by > and _mm512_cmpge_epu64_mask by _mm512_cmpgt_epu64_mask. This approach (with slightly different implementation) has also been suggested for SIMD-based k-way search in (Shahvarani and Jacobsen, 2016; Zhou and Ross, 2002; Schlegel et al., 2009).
@@ 2       $v\leftarrow$ node pointed by entry at position $v[succ_{>k}(v)]$
6else $\triangleright$ $k$ does not exist
@@ 2       $v\leftarrow$ node pointed by entry at position $v[succ_{>k_{1}}(v)]$
5       $v\leftarrow nextleaf(v)$triangleright next to $v$ leaf
@@ 4. Gaps and Updates
The main novelty of BS-tree is the implementation of gaps in nodes using duplicated keys, which facilitates efficient updates and allows the use of Snippet 2 for search at the same time. Specifically, while SIMD-based k-way search on packed arrays has also been suggested before (Shahvarani and Jacobsen, 2016; Zhou and Ross, 2002; Schlegel et al., 2009), to our knowledge it has never been applied on B+-tree nodes with unused slots. As discussed in Sec. 3.1, unused key slots at the end of each node are filled with MAXKEY values; hence, uniqueness is not a requirement for unused key positions. In addition, BS-tree does not require the used key slots to be continuous at the beginning of the node. This means that ‘gaps’ with unused key slots are allowed in a node.

### v2

@@ 1. Introduction
The B+-tree is the dominant indexing method for DBMSs, due to its low and guaranteed cost for (equality and range) query processing and updates. It was originally proposed as a disk-based index, where the objective is to minimize the I/O cost of operations. As memories become larger and cheaper, main-memory and hardware-specific implementations of the B+-tree (BoyarL92; LehmanC86; RaoR99; RaoR00; BohannonMR01; KimCSSNKLBD10; fix2011accelerating; LevandoskiLS13a; Kaczmarski12; ShahvaraniJ16; YanLPZ19; KwonLNNPCM23), as well as alternative access methods for in-memory data (Morrison68; AskitisS10; BohmSVFHL09; KissingerSHL12; MaoKM12; LeisK013; ZhangAPKMS16; BinnaZPSL18; ZhangLLAKKP18) have been proposed. The optimization objective in all these methods is minimizing the computational cost and cache misses during search. More recently, learned indices (KraskaBCDP18; GalakatosMBFK19; FerraginaV20; DingMYWDLZCGKLK20; ZhangG22; WuCYSKX22; KipfMRSKK020; WuZCCWX21; LiLZDYP23; ZhangQYB24), which replace the inner nodes of the B+-tree by ML models have been suggested as a way for reducing the memory footprint of indexing and accelerating search at the same time.
@@
In this paper, we propose BS-tree, a B+-tree for main memory data, which is optimized for modern commodity hardware and data parallelism. BS-tree adopts the structure of the disk-based B+-tree (i.e., a balanced, multiway tree), setting the node size to a memory block that can be processed in parallel. At the heart of our proposal lies a data-parallel successor operator (succsucc), implemented using SIMD, which is applied at each tree level for branching during search and updates and for locating the search key position at the leaf level. To facilitate fast updates, without affecting SIMD-based search, we propose a novel implementation for gaps (unused positions) by duplicating keys. The main idea is that we write in each unused slot the next used key value in the node or a global MAXKEY value if all subsequent slots are unused. Our BS-tree construction algorithm initializes sparse leaf nodes with intentional gaps in them, in order to (i) delay possible splits and (ii) reduce data shifting at insertions. Splitting also adds gaps proactively. Finally, we apply a frame-of-reference (FOR) based compression method that allows nodes that use fixed-size memory blocks to have varying capacities, which saves space and increases data parallelism. The computational cost of search and update operations in BS-tree is $O(\log_{f}n)$, where $f$ is the capacity of the nodes, assuming that $f$ is selected such that each node can be processed by a (small) constant number of SIMD instructions.
@@
Novelty and contributions. There already exist several SIMD-based implementations of B-trees and k-ary search (SchlegelGL09; KimCSSNKLBD10; ShahvaraniJ16; YanLPZ19; KwonLNNPCM23; Graefe24). In addition, updatable learned indices (DingMYWDLZCGKLK20) use gaps to facilitate fast updates. Finally, key compression in B-trees has also been studied in previous work (BohannonMR01). To our knowledge, our proposed BS-tree is the first B+-tree implementation that gracefully combines all these features, achieving at the same time minimal storage and high throughput. In particular, the use of duplicate key vales in unused gaps/slots allows (i) branchless, data-parallel SIMD search at each node, and (ii) efficient key insertions and deletions with limited shifting of keys within each node. In addition, our compression scheme allows for direct application of data-parallel search on compressed nodes. To our knowledge, using duplicate keys in gaps within each node for efficient data-parallel search and updates at the same time is novel and has not been supported by previous B-tree implementations (Graefe24). We implement a version of optimistic lock coupling in BS-tree for concurrency control and extensively compare BS-tree with open-source single- and multi-threaded implementations of state-of-the-art non-learned and learned indices on widely used real datasets, to find that BS-tree and its compressed version consistently prevail in different workloads of reads and updates, typically achieving 1.5x-2x higher throughput than the best competitor.
@@ 2.1. B-tree
The B+-tree is considered the de-facto access method for relational data, having substantial advantages over hash-based indexing with respect to construction cost, support of range queries, sorted data access, and concurrency control (Graefe11; ElmasriN89; Graefe24). As memory sizes grow, the interest has shifted to in-memory access methods (ZhangCOTZ15). Rao and Ross (RaoR99) were the first to consider the impact of cache misses in memory-based data structures; they proposed Cache-Sensitive Search Trees (CSS-trees), in which every node has the same size as the cache-line of the machine and does not need to keep pointers for the links between nodes, but offsets that can be calculated by arithmentic operations. Rao and Ross (RaoR00) also proposed the Cache Sensitive B+-tree (CSB+tree), which achieves cache performance close to CSS-Trees, while having the advantages of a B+-tree. Chen et al. (ChenGM01; ChenGMV02) showed how prefetching can significantly improve the performance of index structures by reducing memory access latency. The pkB-tree (BohannonMR01) is an in-memory variant of the B-tree that uses partial-keys (fixed-size parts of keys), which reduce cache misses and improve search performance. Zhou and Ross (ZhouR03) investigated buffering techniques, based on fixed-size or variable-sized buffers, for memory index structures, aiming to avoid cache thrashing and to improve the performance of bulk lookup in relation to a sequence of single lookups. Graefe and Larson (GraefeL01) surveyed techniques that improve the perfomance of B+-tree by exploting CPU caches.
@@
The advent of SIMD instructions and GPUs opened new perspectives for in-memory index structures. In an early work, Zhou and Ross (ZhouR02) explored the use of SIMD to parallelize key database operations (such as scans, joins, and filtering), minimizing branch mispredictions. Schlegel et al. (SchlegelGL09) present methods for SIMD-based k-ary search (find which out of k partitions contains a search key). FAST (KimCSSNKLBD10) optimizes k-way tree search by leveraging architecture-specific features such as cache locality, SIMD parallelism on CPUs, and massive parallelism on GPUs. (fix2011accelerating) introduced a “braided” B+-tree structure optimized for parallel searches on GPUs, enabling lock-free traversal using additional pointers. Kaczmarski (Kaczmarski12) proposed a bottom-up B+-tree construction and maintenance technique using CPU and GPU for bulk-loading and updates. Bw-Tree (LevandoskiLS13a) is a highly scalable and latch-free B+-tree variant optimized for modern hardware platforms, including multi-core processors and flash storage. Hybrid B+-tree (ShahvaraniJ16) leverages both CPU and GPU resources to optimize in-memory indexing, by dynamically balancing the workload between the CPU and the GPU. Yan et al. (YanLPZ19) proposed a B+-tree tailored for GPU and SIMD architectures. This structure decouples the “key region”, which contains keys of the B+-tree with the “child region”, which is organized as a prefix-sum array and stores only each node’s first child index in the key region. Kwon et al. (KwonLNNPCM23) proposed DB+-tree, a B+-tree with partial keys, that utilizes SIMD and other sequential instructions for fast branching. PALM (SewallCKSD11) is a parallel latch-free variant of the B+-tree, which is optimized for multi-core processors, enabling concurrent search and update operations. Other works explore the implementation of B-trees on flash memory (WuKC07; AgrawalGSDS09; LiH0LY10; JorgensenRSS11; NaLM12; AthanassoulisA14; JinYJYY16; WangZHZ20), non-volatile memory (ChenJ15; LiuCW20) and hardware transactional memory (SiakavarasBNGK20).
@@ 2.3. Other in-memory access methods
Besides B-trees, other data structures have also been used for in-memory indexing, especially trie-based ones, such as the HAT-trie (AskitisS07; AskitisS10), the generalized prefix tree (trie) (BohmSVFHL09), KISS-TREE (KissingerSHL12), and Masstree (MaoKM12). Leis et al. (LeisK013) proposed a fast and space-efficient in-memory trie called ART, which dynamically adjusts its node sizes providing a compact and cache-efficient representation. ART uses lazy expansion and path compression to improve space utilization and search performance. Leis et al. proposed two synchronization protocols for ART in (LeisSK016), which have good scalability despite relying on locks: optimistic lock coupling and the read-optimized write exclusion (ROWEX) protocol. Height Optimized Trie (HOT) (BinnaZPSL18) is an in-memory trie-based index that reduces tree height through path compression and node merging. SuRF (Succinct Range Filter) (ZhangLLAKKP18) leverages succinct tries to provide a space-efficient solution for range query filtering.
@@ 2.4. Learned Indexing
The advent of fast and accurate machine learning techniques inspired the design of a new type of index structure, called learned index (KraskaBCDP18). The main idea is to learn a cumulative distribution function (CDF) of the keys and define a Recursive Model Index (RMI) that replaces the inner nodes of a B+-tree by a hierarchy of models that can predict very fast the position of the search key. FITing tree (GalakatosMBFK19) and PGM-index (FerraginaV20) build upon RMI with a focus on improving model performance. with provable worst-case bounds on query time and space usage. The RadixSpline (RS) (KipfMRSKK020) learned index can be constructed in a single traversal of sorted data. ALEX (DingMYWDLZCGKLK20) is an updatable learned index structure, based on RMI. ALEX utilizes a gapped array layout that gracefully distributes extra space between elements based on the model’s predictions, enabling faster updates and lookups by exponential search. Other updatable learned indices include CARMI (ZhangG22), NFL (WuCYSKX22), LIPP (WuZCCWX21), and DILI (LiLZDYP23), and Hyper (ZhangQYB24). Refs. (MarcusKRSMK0K20; sosd-neurips; WongkhamLLZLW22) provide comprehensive evaluations on updatable learned indices and traditional indices including many important findings, based on tests on several real-world datasets.
@@ 3.1. The structure
BS-tree follows the structure of the B+-tree. Each internal node of the tree fits up to $N$ references to nodes at the lower level and up to N1N-1 keys. Leaf nodes contain rid-key pairs, where a record-id (rid) is the address (potentially on the disk) of the record that has the corresponding key value. We assume that keys are unique (if not, rid is replaced by a pointer to a block that keeps the rid’s of all records having the corresponding search key). The storage of the rid’s is decoupled from the storage of the keys, i.e., they are stored in two different arrays, such that the rid array is accessed only if necessary (i.e., only if the key is found and we need access to the corresponding record). Similarly, the storage of keys in an internal node is decoupled from the storage of children addresses, to facilitate fast search, as we explain later. Each leaf node hosts the address of the next leaf in the total key order; Leaf chaining efficiently supports range queries, where query results are in consecutive tree leaves. For the BS-tree nodes we use a value of $N$ that facilitates fast and parallel search, as we will explain in Sections 5 and 6. For the efficient handling of updates, we allow gaps (i.e., unused slots) in nodes, similarly to previous work (DingMYWDLZCGKLK20; WuZCCWX21; LiLZDYP23), as will be discussed in Section 4.
Figure 1 shows an example of a BS-tree, where each node holds up to N1=4N-1=4 keys. Each non-leaf node is shown as an array of $N$ node pointers (bottom) and N1N-1 keys (top) that work as separators. All keys in the subtree pointed by the $i$-th pointer are strictly smaller than the $i$-th key and greater than or equal to the (i1)(i-1)-th key (if $i>0$). Any unused key slots at the end of each node carry a special MAXKEY value (denoted by $\infty$ in the figure), which is larger than the maximum possible value in the key domain. For example, if keys are unsigned 64-bit integers, MAXKEY $2^{64}-1$ and key values range in $0,2^{64}-2]$.
@@ 3.2. Search within a BS-tree node
1int succG(Node *v, uint64 skey) {
@@
2int succG_SIMD(Node *v, uint64 skey) {
@@
Similarly, $succ_{\geq k}(v)=|\{x:x\in v.keys\land k>x\}|$ and the same code snippets can be used, by replacing comparison >= by > and _mm512_cmpge_epu64_mask by _mm512_cmpgt_epu64_mask. This approach (with slightly different implementation) has also been suggested for SIMD-based k-way search in (ShahvaraniJ16; ZhouR02; SchlegelGL09).
@@ 3  $v\leftarrow$ node pointed by entry at position $v[succ_{>k}(v)]$
7else $\triangleright$ $k$ does not exist
@@ 3  $v\leftarrow$ node pointed by entry at position $v[succ_{>k_{1}}(v)]$
6  $v\leftarrow nextleaf(v)$ $\triangleright$ next to $v$ leaf
@@ 4. Gaps and Updates
The main novelty of BS-tree is the implementation of gaps in nodes using duplicated keys, which facilitates efficient updates and allows the use of Snippet 2 for search at the same time. Specifically, while SIMD-based k-way search on packed arrays has also been suggested before (ShahvaraniJ16; ZhouR02; SchlegelGL09), to our knowledge it has never been applied on B+-tree nodes with unused slots. As discussed in Sec. 3.1, unused key slots at the end of each node are filled with MAXKEY values; hence, uniqueness is not a requirement for unused key positions. In addition, BS-tree does not require the used key slots to be continuous at the beginning of the node. This means that ‘gaps’ with unused key slots are allowed in a node.